4 persons are chosen at random from a group of 3 men, 2 women, 3 children. Find the probability that out of 4 choice, exactly 2 are children.

Grade:Upto college level

4 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Probability that out of 4 exactly 2 children is = (3c2)(5c2)/8c4
=30/70=3/7
Thanks and Regards
Shaik Aasif
askIITians faculty
Saaraansh Mishra
6 years ago
Sir your answer is wrong.It is 10/21.As it was given by a book of fiitjee . So your answer is wrong. Please provide us a valid solution.
Kritika Jain
6 years ago
Can anyone please help me with this one?
Answer is 10/21 and the no. Of children in the group are 4 not 3
Ranganatha
6 years ago
Selecting 2 children is 4C2=6Selecting 2 Non children (3M+2W)= 5C2=10Total ways of selecting 4 persons=9C4=126Probability is (6×10)/126=60/126=10/21